# In this paper, we show that the constant-dimensional Weisfeiler–Leman algorithm for groups (Brachter & Schweitzer, LICS 2020) can be fruitfully used to improve parallel complexity upper bounds on isomorphism testing for several families of groups. In particular, we show:

### v6

@@ 6.1 Preliminaries
Let $S_{1},\dotsc,S_{k}\leq G$ be nonabelian simple subgroups such that for all distinct $i,j\in[k]$ we have $[S_{i},S_{j}]=1$. Then $\langle S_{1},\dotsc,S_{k}\rangle=S_{1}S_{2}\dotsb S_{k}=S_{1}\times\dotsb\times S_{k}$.
@@
By induction on $k$. The base case $k=1$ is vacuously true. Suppose $k\geq 2$ and that the result holds for $k-1$. Then $T:=S_{1}S_{2}\dotsb S_{k-1}=S_{1}\times\dotsb\times S_{k-1}$. Now, since $S_{k}$ commutes with each $S_{i}$, and they generate $T$, we have that $[S_{k},T]=1$. Hence $T$ is contained in the normalizer (or even the centralizer) of $S_{k}$, so $TS_{k}=S_{k}T=\langle T,S_{k}\rangle$, and $S_{k}$ and $T$ are normal subgroups of $TS_{k}$. As $TS_{k}=\langle T,S_{k}\rangle$ and $T,S_{k}$ are both normal subgroups of $TS_{k}$ with $[T,S_{k}]=1$, we have that $TS_{k}$ is a central product of $T$ and $S_{k}$. As $Z(T)=Z(S_{k})=1$, it is their direct product. ∎
@@ 6.2 Groups without Abelian normal subgroups in parallel
It remains to check whether $S\in\text{Fac}(\text{Soc}(G))$. By Cor. 6.6, $S\in\text{Fac}(\text{Soc}(G))$ if and only if $N:=\operatorname{ncl}_{G}(S)=S^{k}$ for some $k$ and $S\unlhd N$. As $S$ is simple, it suffices to check that each conjugate of $S$ is either (1) equal to $S$ or (2) intersects trivially with $S$ and commutes with $S$. For a given $g\in G$ and each $h\in S$, we may check whether $h\in gSg^{-1}$. If there exist non-trivial $h_{1},h_{2}\in S$ such that $h_{1}\in gSg^{-1}$ and $h_{2}\not\in gSg^{-1}$, we return that $S\not\in\text{Fac}(\text{Soc}(G))$. Otherwise, we know that all conjugates of $S$ are either equal to $S$ or intersect $S$ trivially. Next we check that those conjugates that intersect $S$ trivially commute with $S$. For each $g\in G,h_{1},h_{2}\in S$ we check whether $gh_{1}g^{-1}\in S$; if not, we check that $[gh_{1}g^{-1},h_{2}]=1$. If not, then we return that $S\not\in\text{Fac}(\text{Soc}(G))$. If all these tests pass, then $S$ is a direct factor of the socle. For both of these procedures, we only need to iterate over 3- and 4-tuples of elements of $G$ or $S$, so this entire procedure is L-computable. The result follows. ∎
@@
By Prop. 6.8, we may assume that both groups are semisimple. We now note that, by Lem. 6.9, if $\text{Soc}(G)\not\cong\text{Soc}(H)$, then $(O(1),O(1))$-WL Version II will distinguish $G$ from $H$. For in this case, there is some simple normal factor $S\in\text{Fac}(\text{Soc}(G))$ such that there are more copies of $S$ in $\text{Fac}(\text{Soc}(G))$ than in $\text{Fac}(\text{Soc}(H))$. Thus under any bijection Duplicator selects, some element of $S$ must get mapped into a simple direct factor of $\text{Soc}(H)$ that is not isomorphic to $S$, and thus by Lem. 6.9, Spoiler can win with 4 pebbles and 5 rounds.
@@ 7 Count-free Weisfeiler–Leman
We begin by defining analogous pebble games and logics for the three count-free WL versions. Furthermore, we establish the equivalence of the three count-free WL versions up to $O(\log n)$ rounds. These results extend [BS20, Section 3] to the count-free setting.
@@ 7.1 Equivalence between count-free WL, pebble games, and logics
  4. 4.
    The winning condition is checked. This will be formalized later.
Let $v_{1},\ldots,v_{m}$ be the pebbled elements of $G$ at the end of step 1, and let $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ be the corresponding pebbled elements of $H$. Spoiler wins precisely if the map $v_{\ell}\mapsto v_{\ell}^{\prime}$ does not extend to a marked equivalence in the appropriate version of WL. Duplicator wins otherwise. Spoiler wins, by definition, at round 00 if $G$ and $H$ do not have the same number of elements. We will show, for $J\in\{I,II\}$, that $G$ and $H$ are not distinguished by the first $r$ rounds of $k$-WL version $J$ if and only if Duplicator wins the first $r$ rounds of the Version $J$ $(k+1)$-pebble game.
@@
Let $\overline{g}:=(g_{1},\ldots,g_{k})\in G^{k}$ and $\overline{h}:=(h_{1},\ldots,h_{k})\in H^{k}$. If the count-free $(k,r)$-WL distinguishes $\overline{g}$ and $\overline{h}$, then Spoiler can win in the count-free $(k+1)$-pebble game within $r$ moves on the initial configuration $(\overline{g},\overline{h})$. (We use the same version of WL and the pebble game).
@@
  • Version I: For $r=0$, then $\overline{g}$ and $\overline{h}$ differ with respect to the Version I marked equivalence type. Fix $r>0$. Suppose that $\chi_{r}(\overline{g})\neq\chi_{r}(\overline{h})$. We have two cases. Suppose first that $\chi_{r-1}(\overline{g})\neq\chi_{r-1}(\overline{h})$. Then by the inductive hypothesis, Spoiler can win in the $(k+1)$-pebble game using at most $r-1$ moves.
    Suppose instead that $\chi_{r-1}(\overline{g})=\chi_{r-1}(\overline{h})$. So without loss of generality, there exists an $x\in G$ such that the color configuration $(\chi_{r-1}(\overline{g}(g_{1}/x)),\ldots,\chi_{r-1}(\overline{g}(g_{k}/x))$ does not appear amongst the colored $k$-tuples of $H$. Thus, for all $y\in H$, there exists some $j\in[k]$ such that $\chi_{r-1}(\overline{g}(g_{j}/x))\neq\chi_{r-1}(\overline{h}(h_{j}/y))$. Spoiler moves pebble $p_{j}$ to $x$. By the inductive hypothesis, Spoiler wins with $r-1$ additional moves.
@@
Let $\overline{g}:=(g_{1},\ldots,g_{k})\in G^{k}$ and $\overline{h}:=(h_{1},\ldots,h_{k})\in H^{k}$. Suppose that Spoiler can win in the count-free $(k+1)$-pebble game within $r$ moves on the initial configuration $(\overline{g},\overline{h})$. Then the count-free $(k,r)$-WL distinguishes $\overline{g}$ and $\overline{h}$. (We use the same version of WL and the pebble game).
@@
  • Version I: If $r=0$, then the initial configuration is already a winning one for Spoiler. We have in this case, by definition, that $\overline{g},\overline{h}$ receive different colorings at the initial round of count-free $k$-WL. Now let $r>0$, and suppose that Spoiler wins at round $r$ of the pebble game. Suppose Spoiler begins their $r$-round winning strategy by moving pebble $p_{j}$ from $g_{j}$ to $x$, and suppose that Duplicator responds by placing $p_{j}^{\prime}$ on $y$. Thus, Spoiler has a winning strategy in the $(k+1)$-pebble, $(r-1)$-round game, starting from the configuration $(\overline{g}(g_{j}/x),\overline{h}(h_{j}/y))$. So by the inductive hypothesis, $\chi_{k,r-1}(\overline{g}(g_{j}/x))\neq\chi_{k,r-1}(\overline{h}(h_{j}/y))$. By the definition of the count-free WL refinement step, we will thus have that $\chi_{k,r}(\overline{g})\neq\chi_{k,r}(\overline{h})$. The result now follows.
@@ 7.2 Logics
  • Version II: We consider relations of the form $R_{w}(x_{1},\ldots,x_{t})$, where $w\in(\{x_{1},\ldots,x_{t}\}\cup\{x_{1}^{-1},\ldots,x_{t}^{-1}\})^{*}$. We say that $R_{w}(x_{1},\ldots,x_{t})$ is fulfilled by a $t$-tuple $(g_{1},\ldots,g_{t})$ of group elements in $G$ if and only if $w(g_{1},\ldots,g_{t})=1$ in $G$. Define $\mathcal{L}_{II}$ to be the extension of first-order logic obtained by adding all such relations $R_{w}$. Let $\mathcal{L}_{II}^{k}$ be the fragment of $\mathcal{L}_{II}$ that uses at most $k$ variables and relations $R_{w}$, where $w$ ranges through the $k-1$ variable words defined over these $k$ variables and their inverses. We refer to the logic $\mathcal{C}_{II}$ as the logic obtained from $\mathcal{L}_{II}$ by adding counting quantifiers $\exists^{\geq n}x_{i}\,\varphi$ and $\exists{!n}\,\varphi$, and its $k$-variable fragment as $\mathcal{C}_{II}^{k}$.
@@
Brachter & Schweitzer [BS20, Lemma 3.6] showed that for $J\in\{I,II\}$ two $k$-tuples $\overline{g},\overline{h}$ receive a different initial color under $k$-WL Version $J$ if and only if there is a quantifier-free formula in $\mathcal{C}_{J}$ that distinguishes $\overline{g},\overline{h}$. As such formulas do not use any quantifiers, $\overline{g},\overline{h}$ receive a different initial color under $k$-WL Version $J$ if and only if there is a quantifier-free formula in $\mathcal{L}_{J}$ that distinguishes $\overline{g},\overline{h}$. Now the equivalence between the $(k+1)$-pebble, $r$-round Version $J$ count-free pebble game and the $(k+1)$-varible, quantifier-depth $r$ fragment of $\mathcal{L}_{J}$ follows identically from the argument as in the case of graphs [CFI92]. We record this with the following theorem.
@@
Let $G$ and $H$ be groups of order $n$, and let $J\in\{I,II\}$. Fix $k\geq 2$, and let $\overline{g}\in G^{k},\overline{h}\in H^{k}$. Let $\chi_{k,r}$ be the coloring computed by the count-free $(k,r)$-WL on $G$ and $H$. We have that $\chi_{k,r}(\overline{g})\neq\chi_{k,r}(\overline{h})$ if and only if there exists a sentence $\varphi\in\mathcal{L}_{J}$ that uses at most $k+1$ variables and quantifier depth at most $r$, such that $(G,\overline{g})\models\varphi$ and $(H,\overline{h})\not\models\varphi$.
@@ 7.3 Count-free WL and Abelian groups
We now turn to showing that the count-free WL Version II algorithm fails to yield a polynomial-time isomorphism test for even Abelian groups.
@@
A basis of an Abelian group $A$ is a generating set $a_{1},\dotsc,a_{k}$ of $A$ such that $A=\prod_{i=1}^{k}\langle a_{i}\rangle$, or equivalently, $\langle a_{i}\rangle\cap\langle a_{j}:j\neq i\rangle=1$ for all $i\in[k]$. For finite Abelian $p$-groups $A$, all bases have the same size, which is equal to the dimension of $A/A^{p}$ as an $\mathbb{F}_{p}$-vector space. (If $A$ is not a $p$-group this is false, e.g. $\mathbb{Z}/6\mathbb{Z}$ a basis of size 1 and a basis of size 2.) This is also the minimum number of generators of $A$, which is also denoted $d(A)$. If $A$ is an Abelian $p$-group and $x\in A$ is not the $p$-th power of another element of $A$, then there is a subgroup $A^{\prime}\leq A$ such that $A=A^{\prime}\times\langle x\rangle$, or equivalently there is a basis of $A$ that includes $x$ (this follows from Nakayama’s Lemma, by the analogous result for bases of the $\mathbb{F}_{p}$-vector space $A/A^{p}$).
@@
We refer to $(a,b,c)$ as the type of the subgroup $A$. An alternative characterization is: $a$ is the number of direct factors of $A$ that are $\mathbb{Z}/2\mathbb{Z}$ that do not lie in a copy of $\mathbb{Z}/4\mathbb{Z}$ in $G$; $b$ is the the number of direct factors of $A$ that are $\mathbb{Z}/4\mathbb{Z}$; and $c$ is the number of direct factors of $A$ that are $\mathbb{Z}/2\mathbb{Z}$ that are subgroups of a $\mathbb{Z}/4\mathbb{Z}$ in $G$. It is clear that the action of $\operatorname{Aut}(G)$ preserves the type of a subgroup. The converse is also true: by the classification of automorphisms of Abelian groups (see, e.g., [HR07] for a nice exposition) or by using Smith Normal Form as mentioned above, it follows that for any two subgroups $A,B\leq G$ of the same type, there is an automorphism of $G$ that sends $A$ to $B$.
@@
Let $B=\{x_{1},\dotsc,x_{a+b+c}\}$ be a basis of $A$ such that $\{x_{1},\dotsc,x_{a}\}$ generate a subgroup of non-squares-in-$G$ that is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{a}$, $\{x_{a+1},\dotsc,x_{a+b}\}$ is a basis for a subgroup isomorphic to $(\mathbb{Z}/4\mathbb{Z})^{b}$, and $\{x_{a+b+1},\dotsc,x_{a+b+c}\}$ is a basis for a subgroup of square-in-$G$ that is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{c}$. Let $Y$ be a copy of $(\mathbb{Z}/4\mathbb{Z})^{m}$ such that $\{x_{a+1},\dotsc,x_{a+b+c}\}\subseteq Y$, and let $X$ be a copy of $(\mathbb{Z}/2\mathbb{Z})^{n}\leq G$ such that $\{x_{1},\dotsc,x_{a}\}\subseteq X$, $X\cap Y=1$, and $G=\langle X,Y\rangle=X\times Y$.
@@
    If $c>0$, let $x=x_{a+b+1}$; then $x^{2}=1$ and $x$ is a square in $G$, say $y^{2}=x$. If we replace $x_{a+b+1}$ in $B$ with $y$, we get a basis for a subgroup that contains $A$ (since $y^{2}=x_{a+b+1}$), and we have changed the type to $(a,b+1,c-1)$.
@@
    As in the previous part, we can find a $y\in Y$ such that $\langle Y\rangle\cap A=1$, but now we let $B=A\times\langle y^{2}\rangle$, which has type $(a,b,c+1)$. This completes the proof of the lemma. ∎
@@
    On the other hand, if $z^{-1}g^{\prime}_{1}$ is not a square in $G$, then the type of $\langle A_{1},g^{\prime}_{1}\rangle$ is $(a+1,b,c)$. Since this subgroup is generated by at most $n-1$ elements, we have $a+1+b+c\leq n$, and in particular $a+1\leq n$. Since $B_{1}$ has type $(a,b,c)$ and $a<n$, by Lemma 7.5 there is an element of $w$ of order 2 in $H$ that is not a square in $H$ such that $w\notin B_{1}$. We will select Duplicator’s response so that $\varphi$ extends to a good isomorphism that maps $z^{-1}g^{\prime}_{1}$ to $w$. Namely, Duplicator responds with $h^{\prime}_{1}:=\varphi(z)w$. By a similar argument to the preceding paragraph, $\varphi$ extends uniquely to a good isomorphism that maps $g^{\prime}_{1}\mapsto h^{\prime}_{1}$, so the new pebbling is good.

### v7

@@ 6.1 Preliminaries
Let $S_{1},\dotsc,S_{k}\leq G$ be nonabelian simple subgroups such that for all distinct $i,j\in[k]$ we have $S_{i},S_{j}]=1$. Then $\langle S_{1},\dotsc,S_{k}\rangle=S_{1}S_{2}\dotsb S_{k}=S_{1}\times\dotsb\times S_{k}$.
@@
By induction on $k$. The base case $k=1$ is vacuously true. Suppose $k\geq 2$ and that the result holds for k1k-1. Then $T:=S_{1}S_{2}\dotsb S_{k-1}=S_{1}\times\dotsb\times S_{k-1}$. Now, since $S_{k}$ commutes with each $S_{i}$, and they generate $T$, we have that $S_{k},T]=1$. Hence $T$ is contained in the normalizer (or even the centralizer) of $S_{k}$, so $TS_{k}=S_{k}T=\langle T,S_{k}\rangle$, and $S_{k}$ and $T$ are normal subgroups of $TS_{k}$. As $TS_{k}=\langle T,S_{k}\rangle$ and $T,S_{k}$ are both normal subgroups of $TS_{k}$ with $T,S_{k}]=1$, we have that $TS_{k}$ is a central product of $T$ and $S_{k}$. As $Z(T)=Z(S_{k})=1$, it is their direct product. ∎
@@ 6.2 Groups without Abelian normal subgroups in parallel
It remains to check whether $S\in\text{Fac}(\text{Soc}(G))$. By Cor. 6.6, $S\in\text{Fac}(\text{Soc}(G))$ if and only if $N:=\operatorname{ncl}_{G}(S)=S^{k}$ for some $k$ and $S\unlhd N$. As $S$ is simple, it suffices to check that each conjugate of $S$ is either (1) equal to $S$ or (2) intersects trivially with $S$ and commutes with $S$. For a given $g\in G$ and each $h\in S$, we may check whether $h\in gSg^{-1}$. If there exist non-trivial $h_{1},h_{2}\in S$ such that $h_{1}\in gSg^{-1}$ and $h_{2}\not\in gSg^{-1}$, we return that $S\not\in\text{Fac}(\text{Soc}(G))$. Otherwise, we know that all conjugates of $S$ are either equal to $S$ or intersect $S$ trivially. Next we check that those conjugates that intersect $S$ trivially commute with $S$. For each $g\in G,h_{1},h_{2}\in S$ we check whether $gh_{1}g^{-1}\in S$; if not, we check that $gh_{1}g^{-1},h_{2}]=1$. If not, then we return that $S\not\in\text{Fac}(\text{Soc}(G))$. If all these tests pass, then $S$ is a direct factor of the socle. For both of these procedures, we only need to iterate over 3- and 4-tuples of elements of $G$ or $S$, so this entire procedure is L-computable. The result follows. ∎
@@
By Prop. 6.8, we may assume that both groups are semisimple. We now note that, by Lem. 6.9, if $\text{Soc}(G)\not\cong\text{Soc}(H)$, then $(O(1),O(1))$-WL Version II will distinguish $G$ from $H$. For in this case, there is some simple normal factor $S\in\text{Fac}(\text{Soc}(G))$ such that there are more copies of $S$ in $\text{Fac}(\text{Soc}(G))$ than in $\text{Fac}(\text{Soc}(H))$. Thus under any bijection Duplicator selects, some element of $S$ must get mapped into a simple direct factor of $\text{Soc}(H)$ that is not isomorphic to $S$, and thus by Lem. 6.9, Spoiler can win with $O(1)$ pebbles and $O(1)$ rounds.
@@ 7 Count-free Weisfeiler–Leman
We begin by defining analogous pebble games and logics for count-free WL versions I-II. Furthermore, we establish the equivalence of the three count-free WL versions up to $O(\log n)$ rounds. These results extend [BS20, Section 3] to the count-free setting.
@@ 7.1 Equivalence between count-free WL, pebble games, and logics
    The winning condition is checked. This will be formalized later.
  3. 3.
@@
Let $v_{1},\ldots,v_{m}$ be the pebbled elements of $G$ at the end of step 1, and let $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ be the corresponding pebbled elements of $H$. Spoiler wins precisely if the map $v_{\ell}\mapsto v_{\ell}^{\prime}$ does not extend to a marked equivalence in the appropriate version of WL. Duplicator wins otherwise. Spoiler wins, by definition, before the start of round 0 if $G$ and $H$ do not have the same number of elements. We will show, for $J\in\{I,II\}$, that $G$ and $H$ are not distinguished by the first $r$ rounds of the count-free $k$-WL version $J$ if and only if Duplicator wins the first $r$ rounds of the Version $J$ $(k+1)$-pebble game.
@@
Suppose $|G|=|H|$, and let $\overline{g}:=(g_{1},\ldots,g_{k})\in G^{k}$ and $\overline{h}:=(h_{1},\ldots,h_{k})\in H^{k}$. If the count-free $(k,r)$-WL distinguishes $\overline{g}$ and $\overline{h}$, then Spoiler can win in the count-free $(k+1)$-pebble game within $r$ moves on the initial configuration $(\overline{g},\overline{h})$. (We use the same version of WL and the pebble game).
@@
  • Version I: For $r=0$, then $\overline{g}$ and $\overline{h}$ differ with respect to the Version I marked equivalence type. Fix $r>0$. Suppose that $\chi_{r}(\overline{g})\neq\chi_{r}(\overline{h})$. We have two cases. Suppose first that $\chi_{r-1}(\overline{g})\neq\chi_{r-1}(\overline{h})$. Then by the inductive hypothesis, Spoiler can win in the $(k+1)$-pebble game using at most r1r-1 moves.
    Suppose instead that $\chi_{r-1}(\overline{g})=\chi_{r-1}(\overline{h})$. So without loss of generality, there exists an $x\in G$ such that the color configuration $(\chi_{r-1}(\overline{g}(g_{1}/x)),\ldots,\chi_{r-1}(\overline{g}(g_{k}/x))$ does not appear amongst the colored $k$-tuples of $H$. That is, for all $y\in H$, there exists some $j\in[k]$ such that $\chi_{r-1}(\overline{g}(g_{j}/x))\neq\chi_{r-1}(\overline{h}(h_{j}/y))$. Spoiler places an unused pebble on $x$. Duplicator responds by placing the corresponding pebble on some $y\in H$. Let $j$ be such that $\chi_{r-1}(\overline{g}(g_{j}/x))\neq\chi_{r-1}(\overline{h}(h_{j}/y))$. Then at the next round, Spoiler removes the pebble on $p_{j}$ and reuses that. By the inductive hypothesis, Spoiler wins with r1r-1 additional moves.
@@
Suppose $|G|=|H|$, and let $\overline{g}:=(g_{1},\ldots,g_{k})\in G^{k}$ and $\overline{h}:=(h_{1},\ldots,h_{k})\in H^{k}$. Suppose that Spoiler can win in the count-free $(k+1)$-pebble game within $r$ moves on the initial configuration $(\overline{g},\overline{h})$. Then the count-free $(k,r)$-WL distinguishes $\overline{g}$ and $\overline{h}$. (We use the same version of WL and the pebble game).
@@
  • Version I: If $r=0$, then the initial configuration is already a winning one for Spoiler. We have in this case, by definition, that $\overline{g},\overline{h}$ receive different colorings at the initial round of count-free $k$-WL. Now let $r>0$, and suppose that Spoiler wins at round $r$ of the pebble game. Suppose Spoiler begins their $r$-round winning strategy by moving pebble $p_{j}$ from $g_{j}$ to $x$, and suppose that Duplicator responds by placing $p_{j}^{\prime}$ on $y$. Thus, Spoiler has a winning strategy in the $(k+1)$-pebble, (r1)(r-1)-round game, starting from the configuration $(\overline{g}(g_{j}/x),\overline{h}(h_{j}/y))$. So by the inductive hypothesis, $\chi_{k,r-1}(\overline{g}(g_{j}/x))\neq\chi_{k,r-1}(\overline{h}(h_{j}/y))$. As it was a winning strategy for Spoiler to move $p_{j}$ from $g_{j}$ to $x$ at the initial round, we thus have for any $y\in H$ that:
    $(\chi_{k,r-1}(\overline{g}(g_{1}/x)),\ldots,\chi_{k,r-1}(\overline{g}(g_{k}/x)))\neq(\chi_{k,r-1}(\overline{h}(g_{1}/y)),\ldots,\chi_{k,r-1}(\overline{h}(h_{k}/y))).$
    It follows that $\chi_{k,r}(\overline{g})\neq\chi_{k,r}(\overline{h})$, as desired. The result now follows.
@@ 7.2 Logics
  • Version II: We consider relations of the form $R_{w}(x_{1},\ldots,x_{t})$, where $w\in(\{x_{1},\ldots,x_{t}\}\cup\{x_{1}^{-1},\ldots,x_{t}^{-1}\})^{*}$. We say that $R_{w}(x_{1},\ldots,x_{t})$ is fulfilled by a $t$-tuple $(g_{1},\ldots,g_{t})$ of group elements in $G$ if and only if $w(g_{1},\ldots,g_{t})=1$ in $G$. Define $\mathcal{L}_{II}$ to be the extension of first-order logic obtained by adding all such relations $R_{w}$. Let $\mathcal{L}_{II}^{k}$ be the fragment of $\mathcal{L}_{II}$ that uses at most $k$ variables and relations $R_{w}$, where $w$ ranges through the k1k-1 variable words defined over these $k$ variables and their inverses. We refer to the logic $\mathcal{C}_{II}$ as the logic obtained from $\mathcal{L}_{II}$ by adding counting quantifiers $\exists^{\geq n}x_{i}\,\varphi$ and $\exists{!n}\,\varphi$, and its $k$-variable fragment as $\mathcal{C}_{II}^{k}$.
@@
Brachter & Schweitzer [BS20, Lemma 3.6] showed that for $J\in\{I,II\}$ two $k$-tuples $\overline{g},\overline{h}$ receive a different initial color under $k$-WL Version $J$ if and only if there is a quantifier-free formula in $\mathcal{C}_{J}$ that distinguishes $\overline{g},\overline{h}$. As such formulas do not use any quantifiers, $\overline{g},\overline{h}$ receive a different initial color under $k$-WL Version $J$ if and only if there is a quantifier-free formula in $\mathcal{L}_{J}$ that distinguishes $\overline{g},\overline{h}$. Now the equivalence between the $(k+1)$-pebble, $r$-round Version $J$ count-free pebble game and the $(k+1)$-variable, quantifier-depth $r$ fragment of $\mathcal{L}_{J}$ follows identically from the argument as in the case of graphs [CFI92]. We record this with the following theorem.
@@
Let $G$ and $H$ be groups of order $n$, and let $J\in\{I,II\}$. Fix $k\geq 2$, and let $\overline{g}\in G^{k},\overline{h}\in H^{k}$. Let $\chi_{k,r}$ be the coloring computed by the count-free $(k,r)$-WL on $G$ and $H$. We have that $\chi_{k,r}(\overline{g})\neq\chi_{k,r}(\overline{h})$ if and only if there exists a sentence formula $\varphi\in\mathcal{L}_{J}$ that uses at most $k$ free variables and quantifier depth at most $r$, such that $(G,\overline{g})\models\varphi$ and $(H,\overline{h})\not\models\varphi$.
@@ 7.3 Count-free WL and Abelian groups
We now turn to showing that the count-free WL Version II algorithm fails to yield a polynomial-time isomorphism test even for Abelian groups.
@@
A basis of an Abelian group $A$ is a generating set $a_{1},\dotsc,a_{k}$ of $A$ such that $A=\prod_{i=1}^{k}\langle a_{i}\rangle$, or equivalently, $\langle a_{i}\rangle\cap\langle a_{j}:j\neq i\rangle=1$ for all $i\in[k]$. For finite Abelian $p$-groups AA, all bases have the same size, which is equal to the dimension of $A/A^{p}$ as an $\mathbb{F}_{p}$-vector space. (If $A$ is not a $p$-group this is false, e.g. $\mathbb{Z}/6\mathbb{Z}$ a basis of size 1 and a basis of size 2.) This is also the minimum number of generators of $A$, which is also denoted $d(A)$. If $A$ is an Abelian $p$-group and $x\in A$ is not the $p$-th power of another element of $A$, then there is a subgroup $A^{\prime}\leq A$ such that $A=A^{\prime}\times\langle x\rangle$, or equivalently there is a basis of $A$ that includes $x$ (this follows from Nakayama’s Lemma, by the analogous result for bases of the $\mathbb{F}_{p}$-vector space $A/A^{p}$).
@@
We refer to $(a,b,c)$ as the type of the subgroup $A$. An alternative characterization is: $a$ is the number of direct factors of $A$ that are $\mathbb{Z}/2\mathbb{Z}$ that do not lie in a copy of $\mathbb{Z}/4\mathbb{Z}$ in $G$; $b$ is the number of direct factors of $A$ that are $\mathbb{Z}/4\mathbb{Z}$; and $c$ is the number of direct factors of $A$ that are $\mathbb{Z}/2\mathbb{Z}$ that are subgroups of a $\mathbb{Z}/4\mathbb{Z}$ in $G$. It is clear that the action of $\operatorname{Aut}(G)$ preserves the type of a subgroup. The converse is also true: by the classification of automorphisms of Abelian groups (see, e.g., [HR07] for a nice exposition) or by using Smith Normal Form as mentioned above, it follows that for any two subgroups $A,B\leq G$ of the same type, there is an automorphism of $G$ that sends $A$ to $B$.
@@
Let $\beta=\{x_{1},\dotsc,x_{a+b+c}\}$ be a basis of $A$ such that $\{x_{1},\dotsc,x_{a}\}$ generate a subgroup of non-squares-in-$G$ that is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{a}$, $\{x_{a+1},\dotsc,x_{a+b}\}$ is a basis for a subgroup isomorphic to $(\mathbb{Z}/4\mathbb{Z})^{b}$, and $\{x_{a+b+1},\dotsc,x_{a+b+c}\}$ is a basis for a subgroup of square-in-$G$ that is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{c}$. Let $Y$ be a copy of $(\mathbb{Z}/4\mathbb{Z})^{m}$ such that $\{x_{a+1},\dotsc,x_{a+b+c}\}\subseteq Y$, and let $X$ be a copy of $(\mathbb{Z}/2\mathbb{Z})^{n}\leq G$ such that $\{x_{1},\dotsc,x_{a}\}\subseteq X$, $X\cap Y=1$, and $G=\langle X,Y\rangle=X\times Y$.
@@
    If $c>0$, let $x=x_{a+b+1}$; then $x^{2}=1$ and $x$ is a square in $G$, say $y^{2}=x$. If we replace $x_{a+b+1}$ in $\beta$ with $y$, we get a basis for a subgroup that contains $A$ (since $y^{2}=x_{a+b+1}$), and we have changed the type to (a,b+1,c1)(a,b+1,c-1).
@@
    As in the previous part, we can find a $y\in Y$ such that $\langle Y\rangle\cap A=1$, but now we let $B=A\times\langle y^{2}\rangle$, which has type $(a,b,c+1)$.
To see the furthermore, suppose that $A\leq B$ and $d(B)=d(A)+1$. Let $y\in G$ be such that $B=A\cdot\langle y\rangle$. If $y$ is an element of order 4, then we must have $\langle y\rangle\cap A=1$, for otherwise $y^{2}\in A$, there is a basis of $A$ that includes $y^{2}$, and then a basis for $B$ can be gotten from that basis for $A$ by replacing $y^{2}$ by $y$, giving $d(B)=d(A)$, contradicting our assumption. Thus if $y$ has order 4 we are necessarily in case 3.
If $y$ has order 2, then we have $\langle y\rangle=\{1,y\}$, and therefore $\langle y\rangle\cap A=1$ (since $d(B)>d(A)$, so $B$ strictly contains $A$), so we again have $B=A\times\langle y\rangle$. If $y$ is not a square in $G$, then we are in case 2, and if $y$ is a square in $G$ then we are in case 4. This completes the proof of the lemma. ∎
@@
    On the other hand, if $z^{-1}g^{\prime}_{1}$ is not a square in $G$, then the type of $\langle A_{1},g^{\prime}_{1}\rangle$ is $(a+1,b,c)$. Since this subgroup is generated by at most n1n-1 elements, we have $a+1+b+c\leq n$, and in particular $a+1\leq n$. Since $B_{1}$ has type $(a,b,c)$ and $a<n$, by Lemma 7.5 there is an element of $w$ of order 2 in $H$ that is not a square in $H$ such that $w\notin B_{1}$. We will select Duplicator’s response so that $\varphi$ extends to a good isomorphism that maps $z^{-1}g^{\prime}_{1}$ to $w$. Namely, Duplicator responds with $h^{\prime}_{1}:=\varphi(z)w$. By a similar argument to the preceding paragraph, $\varphi$ extends uniquely to a good isomorphism that maps $g^{\prime}_{1}\mapsto h^{\prime}_{1}$, so the new pebbling is good.